# MUSE spectroscopy of the high abundance discrepancy planetary nebula NGC 6153

### v2

@@ 5.5 Oxygen content in the cold region
Following the method described in García-Rojas et al. (2022, see their Eqs. 5 and 6), we compute the mass of oxygen in the form of O+ and O++ contained in the cold region, relative to that contained in the warm region. These ratios depend on the values of the electron temperature and densities in both regions, but not on . The results are presented in Table 5.5, and confirm previous results from Gómez-Llanos & Morisset (2020), R22 and GR22: the amount of oxygen in the rich and cold region is of the same order of magnitude as that embedded in the ”classical” warm nebula.
@@ 6.2 The effect of having two different plasma components in a photoionised nebula
Recently, a detailed photoionisation modeling of NGC 6153 was carried out by Gómez-Llanos & Morisset (2020). These authors found that a given ADF(O2+) could be reproduced by different combinations of “normal” and metal-rich components, each with different abundance contrast factors. This introduces a degeneracy in the ADF-ACF relationship, potentially causing apparently low-ADF PNe to conceal two distinct plasma components and a high ACF. Resolving this degeneracy is crucial to untangle the abundance discrepancy puzzle. Additionally, these authors stated that very high spatial and/or spectral resolution observations were the only means to isolate the contribution of the region from the total Hi emission and resolve this important issue. We have made a preliminary attempt to isolate this contribution by combining our data with that of R22. However, as these previous authors wisely pointed out in their work, reality is complex. Nonetheless, we believe that our work represents a foundational step towards a deeper understanding of this type of PN. The key lies in acquiring data that facilitate reliable measurements of electron temperatures in both the cold and warm components, as well as the temperature of Hi through the Balmer and/or Paschen decrements. These are all necessary observables for computing and, consequently, the ACF.
@@
Based on the values presented in Tab. 5.5, we can calculate the product , which is associated with the gas pressure. The pressure in the warm component seems to be comparable to that in the cold region (only 40% higher). Due to the uncertainties in determining the density of the cold region, we cannot definitively say if the two phases are in pressure equilibrium. This apparent lack of pressure equilibrium is common in this type of object, as the difference in density between both components is not as large as the difference in temperatures. There is at present no explanation of this fact. Could it be due to the fact that the two plasma components might not be in contact, as they are two different ejections of gas at different evolutionary times?
@@ 7 Summary
In this work, we present deep integral-field unit (IFU) spectroscopy of the high ADF (ADF 10) PN NGC 6153. The spectra were obtained with the MUSE spectrograph covering the wavelength range 46009300 Å with effective spectral resolution from R = 1,609 to R = 3,506 for the bluest to the reddest wavelengths, respectively, and with a spatial sampling of of 0.2 arcsec.
We built spatially-resolved maps for 60 emission lines as well as for two continuum regions bluer and redder to the H i Paschen discontinuity. We produced a spatially-resolved map of a neutron-capture emission line, [Kr iv]5867.74, which, owing to the faintness of this line, has only been attempted in one other PN.
@@ Acknowledgements
We want to dedicate this paper to the memory of our dear colleague and friend Claudio Mendoza Guardia (1951-2024) who passed away during the preparation of this paper. His contributions to atomic data computations for nebular purposes have been invaluable to the field. His insistence also made it possible for proper credit to be given to the producers of atomic data in nebular works, as it is usually done in the last years. We sincerely thank Grażyna Stasińska, the referee of this paper, for her valuable comments that enhanced its quality. This paper is based on observations made with ESO Telescopes at the Paranal Observatory under program ID 097.D-0241. VG-LL, JG-R and DJ acknowledge financial support from the Canarian Agency for Research, Innovation and Information Society (ACIISI), of the Canary Islands Government, and the European Regional Development Fund (ERDF), under grant with reference ProID2021010074, from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación (AEI- MCINN) under grants “Espectroscopía de campo integral de regiones H II locales. Modelos para el estudio de regiones H II extragalaácticas” with reference DOI:10.13039/501100011033, and ‘Nebulosas planetarias como clave para comprender la evolución de estrellas binarias” with reference PID-2022-136653NA-I00 (DOI:10.13039/501100011033). We also acknowledge support from project P/308614 financed by funds transferred from the Spanish Ministry of Science, Innovation and Universities, charged to the General State Budgets and with funds transferred from the General Budgets of the Autonomous Community of the Canary Islands by the MCINN. JG-R also acknowledges funds from the AEI-MCINN, under Severo Ochoa Centres of Excellence Programme 2020-2023 (CEX2019-000920-S). DJ also acknowledges support from the AEI-MCINU under grant “Revolucionando el conocimiento de la evolución de estrellas poco masivas” and the the European Union NextGenerationEU/PRTR with reference CNS2023-143910 (DOI:10.13039/501100011033).. CM acknowledges funds from UNAM/DGAPA/PAPIIT IN 101220 and IG 101223. RW acknowledges support from STFC Consolidated grant ST/W000830/1. This work has made use of the computing facilities available at the Laboratory of Computational Astrophysics of the Universidade Federal de Itajubá (LAC-UNIFEI). The LAC-UNIFEI is maintained with grants from CAPES, CNPq and FAPEMIG.
@@
The raw MUSE data are available from the ESO archive facility at http://archive.eso.org/. The data products used in this paper: reduced data cubes, extracted emission line maps, analysis pipeline and its outputs, are available from the authors under reasonable request. The emission line maps and the python scripts used for the analysis and to produce the tables and figures presented in this paper are available at https://github.com/VGomezLlanos/NGC-MUSE.
@@ Appendix A He+/H+ abundance maps
Recipe 1: The most simple case is to adopt a general ionisation scheme with 2 zones, where [N ii]5755/6548[S ii]6716/6731 are the physical conditions for ions with IP¡17 eV, and [S iii]6312/9069[Cl iii]5518/5538 for ions with IP17 eV. These conditions are used to compute the emissivities from CELs, ORLs, and the H line.
@@
Recipe 3: We follow a similar scheme as in Recipe 2, but we use the Pashen jump temperature to derive the H emissivity in the determination of the ionic abundances, by setting the
@@ Appendix C Emission line flux maps and line fluxes table for the integrated spectra
In this section we show the derredened flux maps for emission lines of 8 elements heavier than H at different ionisation stages, as wel as the table with the observed and derredened fluxes of 60 emission lines in the integrated spectra of NGC 6153.
![Refer to caption](x26.png)
a Derredened fluxes corrected from recomb. contribution

### v3

@@ 5.5 Oxygen content in the cold region
Following the method described in García-Rojas et al. (2022, see their Eqs. 5 and 6), we compute the mass of oxygen in the form of O+ and O++ contained in the cold region, relative to that contained in the warm region. These ratios depend on the values of the electron temperature and densities in both regions, but not on . The results are presented in Table 8, and confirm previous results from Gómez-Llanos & Morisset (2020), R22 and GR22: the amount of oxygen in the rich and cold region is of the same order of magnitude as that embedded in the “classical” warm nebula.
NGC 6153 | |
[K] | 8,300 |
[K] | 2,000 |
[cm-3] | 3,400 |
[cm-3] | 10,000 |
Mc/Mw(O+) | 1.2 |
Mc/Mw(O2+) | 0.7 |
@@ 6.2 The effect of having two different plasma components in a photoionised nebula
Recently, a detailed photoionisation modelling of NGC 6153 was carried out by Gómez-Llanos & Morisset (2020). These authors found that a given ADF(O2+) could be reproduced by different combinations of “normal” and metal-rich components, each with different abundance contrast factors. This introduces a degeneracy in the ADF-ACF relationship, potentially causing apparently low-ADF PNe to conceal two distinct plasma components and a high ACF. Resolving this degeneracy is crucial to untangle the abundance discrepancy puzzle. Additionally, these authors stated that very high spatial and/or spectral resolution observations were the only means to isolate the contribution of the region from the total Hi emission and resolve this important issue. We have made a preliminary attempt to isolate this contribution by combining our data with that of R22. However, as these previous authors wisely pointed out in their work, reality is complex. Nonetheless, we believe that our work represents a foundational step towards a deeper understanding of this type of PN. The key lies in acquiring data that facilitate reliable measurements of electron temperatures in both the cold and warm components, as well as the temperature of Hi through the Balmer and/or Paschen decrements. These are all necessary observables for computing and, consequently, the ACF.
@@
Based on the values presented in Tab. 8, we can calculate the product , which is associated with the gas pressure. The pressure in the warm component seems to be comparable to that in the cold region (only 40% higher). Due to the uncertainties in determining the density of the cold region, we cannot definitively say if the two phases are in pressure equilibrium. This apparent lack of pressure equilibrium is common in this type of object, as the difference in density between both components is not as large as the difference in temperatures. There is at present no explanation of this fact. It is possible that the two plasma components might not be in contact because they originate from two different ejections of gas at different evolutionary times.
@@ 7 Summary
In this work, we present deep integral-field unit (IFU) spectroscopy of the high ADF (ADF 10) PN NGC 6153. The spectra were obtained with the MUSE spectrograph covering the wavelength range 46009300 Å with effective spectral resolution from R = 1,609 to R = 3,506 for the bluest to the reddest wavelengths, respectively, and with a spatial sampling of 0.2 arcsec.
We built spatially resolved maps for 60 emission lines, as well as for two continuum regions bluer and redder to the H i Paschen discontinuity. We produced a spatially resolved map of a neutron-capture emission line, [Kr iv]5867.74, which, owing to the faintness of this line, has only been attempted in one other PN.
@@ Acknowledgements
We want to dedicate this paper to the memory of our dear colleague and friend Claudio Mendoza Guardia (1951-2024) who passed away during the preparation of this paper. His contributions to atomic data computations for nebular purposes have been invaluable to the field. His insistence also made it possible for proper credit to be given to the producers of atomic data in nebular works, as it is usually done in the last years. We sincerely thank Grażyna Stasińska, the referee of this paper, for her valuable comments that enhanced its quality. This paper is based on observations made with ESO Telescopes at the Paranal Observatory under program ID 097.D-0241. VG-LL, JG-R and DJ acknowledge financial support from the Canarian Agency for Research, Innovation, and Information Society (ACIISI), of the Canary Islands Government, and the European Regional Development Fund (ERDF), under grant with reference ProID2021010074, from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación (AEI- MCINN) under grants “Espectroscopía de campo integral de regiones H II locales. Modelos para el estudio de regiones H II extragalaácticas” with reference DOI:10.13039/501100011033, and “Nebulosas planetarias como clave para comprender la evolución de estrellas binarias” with reference PID-2022-136653NA-I00 (DOI:10.13039/501100011033). We also acknowledge support from project P/308614 financed by funds transferred from the Spanish Ministry of Science, Innovation and Universities, charged to the General State Budgets and with funds transferred from the General Budgets of the Autonomous Community of the Canary Islands by the MCINN. JG-R also acknowledges funds from the AEI-MCINN, under Severo Ochoa Centres of Excellence Programme 2020-2023 (CEX2019-000920-S). DJ also acknowledges support from the AEI-MCINU under grant “Revolucionando el conocimiento de la evolución de estrellas poco masivas” and the European Union NextGenerationEU/PRTR with reference CNS2023-143910 (DOI:10.13039/501100011033). CM acknowledges funds from UNAM/DGAPA/PAPIIT IN 101220 and IG 101223. RW acknowledges support from STFC Consolidated grant ST/W000830/1. This work has made use of the computing facilities available at the Laboratory of Computational Astrophysics of the Universidade Federal de Itajubá (LAC-UNIFEI). The LAC-UNIFEI is maintained with grants from CAPES, CNPq and FAPEMIG.
@@
The data products used in this paper: raw and reduced MUSE data cubes and extracted emission line maps, are available from the ESO archive facility at http://archive.eso.org/. The analysis pipeline, the emission line maps and the python scripts used for the analysis and to produce the tables and figures presented in this paper are available at https://github.com/VGomezLlanos/NGC-MUSE.
@@ Appendix A He+/H+ abundance maps
Recipe 1: The most simple case is to adopt a general ionisation scheme with 2 zones, where [N ii]5755/6548[S ii]6716/6731 are the physical conditions for ions with IP¡17 eV, and [S iii]6312/9069[Cl iii]5518/5538 for ions with IP17 eV. These conditions are used to compute the emissivities from CELs, RLs, and the H line.
@@
Recipe 3: We follow a similar scheme as in Recipe 2, but we use the Paschen jump temperature to derive the H emissivity in the determination of the ionic abundances, by setting the
@@ Appendix C Emission line flux maps and line fluxes table for the integrated spectra
In this section, we show the derredened flux maps for emission lines of 8 elements heavier than H at different ionisation stages, as well as the table with the observed and derredened fluxes of 60 emission lines in the integrated spectra of NGC 6153.
Line | (Å) | F() | I() | Notes |
O ii | 4641.81 | 3.6 0.2 | 4.3 0.3 | blend with N iii 4641.85 |
O ii | 4649.13 | 1.85 0.09 | 2.18 0.16 | blend with O ii 4650.85 |
[Fe iii] | 4658.17 | 0.068 0.004 | 0.080 0.006 | |
O ii | 4661.63 | 0.42 0.02 | 0.49 0.04 | |
He ii | 4685.71 | 11.6 0.6 | 13.4 0.9 | |
[Ar iv] | 4711.37 | 1.5 0.1 | 1.7 0.2 | |
He i | 4713.14 | 1.27 0.09 | 1.42 0.13 | |
[Ne iv] | 4724.15 | 0.022 0.001 | 0.025 0.002 | blend with [Ne iv] 4625.85 |
[Ar iv] | 4740.17 | 2.2 0.1 | 2.4 0.2 | |
N ii | 4802.79 | 0.13 0.01 | 0.14 0.01 | |
H i | 4861.32 | 100.0 5.0 | 100.0 5.0 | |
He i | 4921.93 | 1.82 0.09 | 1.73 0.12 | |
[O iii] | 4958.91 | 316.5 15.8 | 293.2 20.1 | |
[Ar iii] | 5191.82 | 0.088 0.004 | 0.069 0.005 | |
[N i] | 5197.90 | 0.23 0.01 | 0.17 0.01 | |
[N i] | 5200.26 | 0.055 0.003 | 0.042 0.003 | |
C ii | 5342.38 | 0.155 0.008 | 0.108 0.007 | |
[Kr iv] | 5346.02 | 0.029 0.002 | 0.020 0.001 | |
He ii | 5411.52 | 1.51 0.08 | 1.01 0.06 | blend with [Fe iii] 5412.00 |
[Cl iii] | 5517.71 | 0.89 0.04 | 0.56 0.03 | |
[Cl iii] | 5537.88 | 1.07 0.05 | 0.66 0.04 | |
N ii | 5666.63 | 0.36 0.02 | 0.21 0.01 | |
N ii | 5676.02 | 0.164 0.008 | 0.094 0.006 | |
N ii | 5679.56 | 0.83 0.04 | 0.47 0.03 | |
N ii | 5686.21 | 0.123 0.006 | 0.070 0.004 | |
N ii | 5710.77 | 0.126 0.006 | 0.071 0.004 | |
[N ii] | 5754.64 | 1.55 0.08 | 0.51 0.05a | |
[Kr iv] | 5867.74 | 0.097 0.005 | 0.050 0.003 | |
He i | 5875.66 | 36.6 1.8 | 18.8 1.1 | |
[O i] | 6300.30 | 1.96 0.10 | 0.82 0.05 | |
[S iii] | 6312.10 | 3.3 0.2 | 1.4 0.1 | |
[O i] | 6363.78 | 0.64 0.03 | 0.26 0.02 | |
C ii | 6461.95 | 0.59 0.03 | 0.23 0.01 | |
[N ii] | 6548.04 | 39.9 2.0 | 15.0 0.9 | |
H i | 6562.80 | 735.1 36.8 | 274.4 14.6 | |
C ii | 6578.05 | 1.64 0.08 | 0.61 0.04 | |
[N ii] | 6583.45 | 125.1 6.3 | 46.3 2.7 | |
He i | 6678.15 | 14.2 0.7 | 5.1 0.3 | |
[S ii] | 6716.44 | 9.6 0.5 | 3.4 0.2 | |
[S ii] | 6730.82 | 15.9 0.8 | 5.5 0.3 | |
[Ar v] | 7005.40 | 0.068 0.003 | 0.021 0.001 | |
He i | 7065.22 | 15.7 0.8 | 4.8 0.3 | |
[Ar iii] | 7135.80 | 71.1 3.6 | 21.2 1.3 | |
[Ar iv] | 7170.50 | 0.223 0.011 | 0.066 0.004 | |
He ii | 7177.52 | 0.485 0.024 | 0.142 0.008 | |
C ii | 7231.33 | 3.3 0.2 | 1.0 0.1 | blend with C ii 7236.42,7237.17 |
[Ar iv] | 7262.70 | 0.152 0.008 | 0.043 0.003 | |
He i | 7281.35 | 2.5 0.1 | 0.7 0.0 | |
[O ii] | 7318.92 | 7.8 0.4 | 1.1 0.1a | blend with [O ii] 7319.99 |
[O ii] | 7329.67 | 6.3 0.3 | 0.9 0.1a | blend with [O ii] 7330.73 |
[Cl iv] | 7530.54 | 1.00 0.05 | 0.26 0.02 | |
[Ar iii] | 7751.10 | 20.7 1.0 | 5.0 0.3 | |
O i | 7771.94 | 0.57 0.04 | 0.14 0.01 | blend with O i 7774.17,7775.39 |
[Cl iv] | 8045.62 | 2.6 0.1 | 0.6 0.0 | |
H i | 8100 | 0.276 0.014 | 0.061 0.004 | Blue continuum PJ |
C iii | 8196.48 | 0.589 0.029 | 0.126 0.008 | |
H i | 8400 | 0.114 0.006 | 0.023 0.001 | Red continuum PJ |
H i | 8750.47 | 6.1 0.3 | 1.1 0.1 | |
H i | 8862.78 | 7.8 0.4 | 1.4 0.1 | |
H i | 9014.91 | 9.0 0.5 | 1.6 0.1 | |
[S iii] | 9068.60 | 202.7 10.1 | 35.3 2.3 | |
H i | 9229.01 | 15.4 0.8 | 2.6 0.1 |
a Derredened fluxes corrected from recombination contribution